How much energy is required to raise the temperature of 1.80 kg of cadmium from 20.7°C to 126°C?
Answer #1

Given,

Mass of Cadmium, m = 1.80 kg

Specific heat of Cadmium, c = 0.23 kJ/Kg.C

Initial temperature, T = 20.7 C

Final temperature, T' = 126 C

We know that heat required,

=> Q = m*c*(T' - T)

=> Q = 1.80*0.23*(126-20.7)

          = 1.80*0.23*105.3

          = 43.6 kJ